OpenStudy (gabylovesyou):

ho do i fix 3 over sqrt(2)

OpenStudy (dumbcow):

is it broken ? :)

OpenStudy (luigi0210):

Haha

OpenStudy (goformit100):

@Gabylovesyou Madam please elaborate your question

OpenStudy (whpalmer4):

To "fix" it, multiply by \[\frac{\sqrt{2}}{\sqrt{2}}\]and simplify.

OpenStudy (whpalmer4):

To "fix" more complicated expressions, such as \[\frac{1}{a-b}\]try multiplying by the conjugate of the denominator: \[\frac{1}{a-b}*\frac{a+b}{a+b} = \frac{a+b}{(a+b)(a-b)} = \frac{a+b}{a^2-b^2}\] Notice that if \(a\) or \(b\) had contained a square root, we would have eliminated the radical sign from the denominator by doing this.